solve: /3x^2+2x/=x/3x+2/


1
Expert's answer
2021-12-09T13:37:28-0500

"3x^2+2x=\frac{x}{3x+2}\\
9x^3+6x^2+6x^2+4x=x\\
9x^3+12x^2+3x=0\\
3x(3x^2+4x+1)=0\\
x(3x+1)(x+1)=0\\
x=0,-1,\frac{-1}{3}"